OpenStudy (anonymous):

two guys are 12 and 32 years old respectively. How much time do is needed for the older one be twice the age of the younger?

OpenStudy (anonymous):

\[\frac{32+t}{12+t}=2\]

OpenStudy (anonymous):

24+2t=32+t

OpenStudy (anonymous):

I think the answer is 8

OpenStudy (anonymous):

Yes 32+8=40 12+8=20

OpenStudy (anonymous):

Sorry I was wrong

OpenStudy (anonymous):

you are right it is 8, i think

OpenStudy (anonymous):

yeah, confirmed